Testbank, Question 50

The stock of Townships Ski Resorts Inc. just paid a dividend of $0.78. What is the expected capital gains yield if the stock is selling for $28.25 today and the required rate of return is 15 percent?

13.62%

12.30%

11.91%

10.73%
Answer #1

28.25 = 28.25*(0.15-g) = 4.2375-28.25g 4.2375-0.78 3.4575 = 29.03g 0.78*(1+g)/(0.15-g) 0.78+0.78g 0.78+0.78g 0.78g+28.25g 0.7